In a constant-pressure calorimeter, 60.0 mL of 0.300 M Ba(OH)2 was added to 60.0 mL of 0.600 M HCl. The reaction caused the temperature of the solution to rise from 24.37 °C to 28.46 °C. If the solution has the same density and specific heat as water, what is ΔH for this reaction (per mole of H2O produced)? Assume that the total volume is the sum of the individual volumes. Find kJ/mol of H20.
Answer #1

60.0 mL of 0.300 M Ba(OH)2 was added to 60.0 mL of 0.600 M HCl.

Moles of Ba(OH)2 added = 0.300 * 0.06

= 0.018

Mass of Ba(OH)2 = 0.018 moles * 171.34 g/mol

= 3.084 g

Moles of HCl added = 0.060 * 0.60

= 0.036

Mass of HCl = 0.036 moles * 36.46 g/mol

= 1.312 g

Total mass of mixture = 3.084 + 1.312

= 4.396 g

Rise in temperature = 28.46 - 24.37 = 4.09 oC

Specific heat = 4.184 J g-1.oC-1

Heat released, q = mC\DeltaT

= 4.396 * 4.184 * 4.09

= 75.228 J

Ba(OH)2 + 2HCl   \rightarrow   BaCl2 + 2H2O

Moles of H2O produced = 0.036

\DeltaHrxn (per mole of H2O produced) = q / m

= 75.228 / 0.036

\DeltaHrxn (per mole of H2O produced) = 2.089 kJ/mol